# Efficient Frame Extraction: A Novel Approach Through Frame Similarity and Surgical Tool Tracking for Video Segmentation

### v1

@@ VII State-of-the-Art Methods
In terms of key frames recognition, our proposed method differs from a recent paper [62], although both aim to detect key frames within a video. First, our approach uses YOLOv8 to detect the presence of these tools in the scene and track their locations across frames. In contrast, the other approach relies on color, motion, and texture and uses a less fine-grained method for motion detection, tracking all moving objects without focusing on specific tools. Second, our goal is to use key frames for phase segmentation, which requires more frames than the other approach, whose objective is merely to select representative frames for summarizing the video, requiring fewer frames. Additionally, we use the X3D CNN model to analyze the temporal information of 2-D frame sequences instead of the hidden Markov model.
Previous research conducted on the PJ dataset using the X3D CNN model employed the Uniform Frame Sampling method (6 fps) [3], which is neither adaptive nor explainable. In this study, we build upon that work by introducing a new method, Kinematics Adaptive Frame Recognition, to selectively capture key frames rather than using all frames. This approach reduces the number of frames by fivefold and improves accuracy by approximately $2\%$. In single-task operating systems, this reduction proportionally decreases the time required for training the model. Theoretically, in real multitasking operating systems, where multiple tasks can run simultaneously, the time savings align closely with the reduction in data size. However, in near-multitasking operating systems, such as Windows (R) operating systems, the time reduction may not scale linearly because the main processes are influenced by competing user and system processes, leading to potential bottlenecks. In our experiments, the runtime required for one epoch of the total training data was approximately 19’15”, while for $10\%$ of the data, it was roughly 5’24”. Although the computation time was reduced, it did not scale directly with the reduction in data size. While an in-depth analysis of the effectiveness of multitasking computing systems falls beyond the scope of this research, it is important to note that reducing the number of frames would not only significantly decrease the storage requirements but also improve computational efficiency, making the method more resource-efficient and practical for large-scale applications.
Last but not least, traditional compression techniques, such as those used by the Moving Picture Experts Group (MPEG) standard [63], primarily focus on reducing file size by leveraging spatial and temporal redundancies in video data, without considering the semantic relevance of specific content. While effective for general-purpose data compression, these methods fail to account for the nuanced actions performed by surgeons, which are critical for surgical analysis. In contrast, our proposed method not only reduces the dataset size but also preserves the semantic significance of surgical actions by tracking the movements of tools. By incorporating this context-aware approach, it achieves improved performance alongside data size reduction. This dual benefit highlights the advantage of the proposed method for specialized applications like surgical skill assessment.
VIII Discussion
In surgery, operative videos are often hierarchically segmented into phases, steps and tasks to describe the procedure at various levels of detail [64]. Each segment is typically defined by a start time and an end time and often spans a considerable duration (from minutes to hours). Often, many portions of the procedure may be idle, for instance, when tools are exchanged or when the camera is cleared of condensation. Moreover, surgical tools are critical to identify in each phase as they are used for performing the surgery, including for retraction, resection, blunt and sharp dissection, cauterizing bleeding and irrigating the surgical field. In our work, by tracking the surgical tools and using KAFR to adaptively select key frames for phase detection resulted in improvements that exceeded our expectations. Initially, we anticipated that using KAFR would involve a trade-off between dataset size and performance. In other words, we expected the accuracy to decrease to some extent for reduction of data size [65, 66]. However, the results indicate otherwise: accuracy can improve even as the dataset size decreases. As demonstrated in the previous section, when the left-hand tool moved slowly, fewer frames were selected. The removed frames had little effect on recognizing the actions [67, 2] performed by the surgeon. Since the model was trained using a frame-by-frame approach, the removed frames introduced confusion during training. Eliminating these frames ultimately resulted in the improvement.
One potential issue with our approach is that using YOLO to track surgical tools often results in loss of tracking either due to tools being retracted or occluded. When the tool is visible or tracked again, a new tracking number is automatically assigned to it, as if it were an instance of new object, needing to continuously monitor the tracking id and assign them to the appropriate tools during post processing. In our experiment, left-hand and right-hand tools are detected based on their centroids relative to a vertical line at the center of the screen. If we track only one tool on either side of the screen, it may move from one side to the other, which can lead to loss of tracking. However, since these tools tend to stay near the center of the screen–convenient for surgeons to focus on–the right-hand or left-hand tool typically remains on its respective side, minimizing the effects. Additionally, because we track tools on both the left and right sides of the screen, the system continues tracking even when a tool crosses to the opposite side. To mitigate this issue, identical tools could be painted with different colors or labeled with unique numbers. However, as our videos were collected for a retrospective study, we were unable to implement or evaluate this setting. This limitation highlights an area for potential future improvement.
Another limitation is that our method assumes the endoscopic camera remains stationary. Changes in the camera view (zoom in and out or excessive movements) can affect the computation of distances, reducing the accuracy of the method. Frequent camera movement similarly impacts other AI models, making this a broader challenge in the field. Our datasets were obtained from robotic procedures where the camera movements are limited compared to laparoscopic procedures where it is controlled manually.
Finally, our methods (Adaptive 1 and Adaptive 2) assume the presence of surgical tools. While we tested it on two datasets, GJ and PJ, and it can be applied to a broader range of datasets, the approach remains restricted to surgical fields. In other domains, the method may face challenges if specific objects are unavailable for tracking.
IX Conclusion
@@
In this study, we introduced Kinematics Adaptive Frame Recognition, a novel method aimed at reducing dataset size and computation time by removing redundant frames while preserving critical information to improve accuracy. The KAFR approach consists of three primary phases: i) Use a YOLOv8 model to detect surgical tools in the scene; ii) Compute similarities between consecutive frames by analyzing spatial positions and velocities of detected tools; iii) Train a X3D CNN to perform phase segmentation based on extracted frames.
We evaluated the efficacy of KAFR on both the GJ and PJ dataset over a span of 10–12 weeks. By adaptively selecting relevant frames, we achieved a tenfold reduction in number of frames while improving accuracy by more than $4\%$ on GJ dataset and fivefold reduction in frames with an approximately $2\%$ accuracy improvement on the PJ dataset. Furthermore, we conducted a comparative analysis with the state-of-the-art method MSE, to demonstrate the competitive performance and efficiency of our approach. Moreover, KAFR can be used to enhance existing methods, providing a valuable improvement to the performance of other AI models. In addition to boosting performance, extracting key frames from a video enhances interpretability and explainability. This process simplifies complex information, clarifies the rationale behind decisions, and links outputs to specific moments. Overall, it makes video-based systems more user-friendly and trustworthy.

### v3

@@ VII Comparison with the state of the art
In terms of key frames recognition, our proposed method differs from a recent paper [..656565removed: in the field ] [loukas2018keyframe], although both aim to detect key frames within a video. First, our approach [..666666removed: identifies significant frames based on the movement of a limited set of tools, including their velocity and acceleration. Our model ]uses YOLOv8 to detect the presence of these tools in the scene and track their locations across frames. In contrast, the other approach relies on color, motion, and texture and uses a less fine-grained method for motion detection, tracking all moving objects without focusing on specific tools. Second, our goal is to use key frames for phase segmentation or gesture detection, which requires more frames than the other approach, whose objective is merely to select representative frames for summarizing the video, requiring fewer frames. Additionally, we use the X3D CNN model to analyze the temporal information of 2-D frame sequences instead of the HMMAR model.
VIII Conclusion
@@
In this study, we introduced Kinematics Adaptive Frame Recognition (KAFR), a novel method aimed at reducing dataset size and computation time by removing redundant frames while preserving critical information to improve accuracy. The KAFR approach consists of three primary phases: i) Use a YOLOv8 model to detect surgical tools in the scene; ii) Compute similarities between consecutive frames by analyzing spatial positions and velocities of detected tools; iii) Train a X3D CNN to perform phase segmentation based on extracted frames.
We evaluated the efficacy of KAFR on [..676767removed: a GJ ] both the GJ and PJ dataset over a span of [..686868removed: 6â ] €“ [..696969removed: 8 ] 11 weeks. By adaptively selecting relevant frames, we achieved a tenfold reduction in number of [..707070removed: frame ] frames while improving accuracy by more than 3% on GJ dataset and fivefold reduction in frames with an approximately 2% accuracy improvement on the PJ dataset. Furthermore, we conducted a comparative analysis with the state-of-the-art method MSE, to demonstrate the [..717171removed: superior ] competitive performance and efficiency of our approach. Moreover, KAFR can be used to enhance existing methods, providing a valuable improvement to the performance of other AI models.
@@ X Data Availability
[Uncaptioned image] Huu Phong Nguyen received a BSc in Physics from Vietnam National University, Hanoi, an MSc in Information Technology from Shinawatra University, Thailand, and a PhD degree from the University of Coimbra, Portugal. He was a member of the Adaptive Computation group at the Centre of Informatics and Systems at the University of Coimbra (CISUC). Currently, he is pursuing a PostDoc at the University of Texas Medical Center, USA. His research interests include medical imaging, image classification, action recognition, pattern recognition, and deep learning.
[Uncaptioned image] Shekhar Madhav Khairnar is currently working as a Data Scientist at UT Southwestern Medical Center, focusing on developing AI models, particularly in computer vision domain. He earned a Bachelor of Engineering (BE) in Electrical from the University of Pune, India, and a Master of Science (MS) in Industrial Engineering from the University of Texas at Arlington, where he received awards including the GTA/GRA and G.T. Stevens Endowed Scholarship. His research areas include computer vision, covering image classification, segmentation, object detection, and tracking, and broader fields such as machine learning, natural language processing (NLP), large language models (LLMs), and deep learning.
[Uncaptioned image] Sofia Garces Palacios is a Research Associate at the Artificial Intelligent and Medical Simulation (AIMS) Laboratory within the Department of Surgery at UT Southwestern Medical Center. Dr. Sofia holds a Doctor of Medicine degree from the University of Manizales in Colombia and is pursuing a career in Surgery. Her research interests are centered on surgical education, minimally invasive surgical techniques, and global surgery.
[Uncaptioned image] Amr Al-Abbas is a General Surgery Resident at UT Southwestern Medical Center and Parkland Hospital, with over five years of experience. He also served as a Postdoctoral Research Fellow at UT Southwestern, focusing on artificial intelligence for video-based assessment in robotic pancreatic surgery. Previously, Dr. Al Abbas completed postdoctoral research in surgical oncology at UPMC and the University of Pittsburgh, where he conducted clinical outcomes research. He has additional research experience in bariatric surgery at the American University of Beirut. His academic background includes work in cell and molecular biology and surgical oncology sub-internships at Duke University and the American University of Beirut. Dr. Amr Al Abbas is interested in surgical oncology robotic, surgery surgical, education video-based assessment, machine learning.
[Uncaptioned image] Melissa E. Hogg is the Director of HPB Surgery at Northshore University Hospital and Director of Advanced Robotic Training for the Grainger Center for Simulation and Innovation. She did her medical school and residency training at Northwestern University in Chicago, IL, and her fellowship and masters training in Medical Education and Clinical Research at the University of Pittsburgh, PA where she also had her first faculty appointment. She is faculty of the University of Chicago Pritzker School of Medicine and serves as Associate Program Director of the General Surgery Residency.
[Uncaptioned image] Amer H. Zureikat is chief of the Division of Gastrointestinal (GI) Surgical Oncology at UPMC Hillman Cancer Center, co-director of the UPMC Hillman Cancer Center Pancreatic Cancer Program, and associate professor of surgery at the University of Pittsburgh School of Medicine. Dr. Zureikat specializes in cancers and diseases of the pancreas, stomach, liver, and duodenum, and practices state-of-the-art robotic surgery.
[Uncaptioned image] Patricio Polanco is an Associate Professor of Surgical Oncology at UT Southwestern Medical Center, specializing in minimally invasive and robotic surgery for pancreatic and liver cancers. He directs the Robotic Surgery Training Program and co-directs the Pancreatic Cancer Program and Prevention Clinic. Dr. Polanco leads the Peritoneal Surface Malignancies and HIPEC Program at the Harold C. Simmons Comprehensive Cancer Center. Originally from Peru, he completed his medical education at Universidad de San Martin de Porres and further specialized in complex surgical oncology in the U.S. He is double board-certified and has published over 100 peer-reviewed articles, earning numerous research awards. Dr. Polanco is a recognized leader in robotic surgery training and has served in various leadership roles within national and international surgical associations. His contributions to global health include training surgeons and participating in medical missions across Latin America.
[Uncaptioned image] Herbert J. Zeh III is Professor and Chair of the Department of Surgery at UT Southwestern Medical Center, where he holds the Hall and Mary Lucile Shannon Distinguished Chair in Surgery. A leader in pancreatic diseases and cancer, Dr. Zeh has over 20 years of surgical experience. He earned his medical degree from the University of Pittsburgh and completed his residency and fellowship at The Johns Hopkins Hospital. Before joining UT Southwestern in 2018, he served as the Watson Family Professor of Surgery at the University of Pittsburgh and co-directed the UPMC Pancreatic Cancer Center. A prolific author with over 200 peer-reviewed publications, Dr. Zeh’s research focuses on novel treatments for pancreatic cancer and robotic surgery innovations. He is a member of numerous prestigious surgical organizations.
[Uncaptioned image] Ganesh Sankaranarayanan received the BE degree in electrical electronics engineering from Annamalai University in 1997, the MS degree in electrical engineering from the University of Texas at Arlington in 2002, and the PhD degree in electrical engineering from the University of Washington in 2007. He was a postdoctoral research associate from 2008-2010 and research assistant professor from 2010-2013 in the Department of Mechanical, Aerospace, and Nuclear Engineering. He was a research scientist at the Center for Modeling, Simulation, and Imaging in Medicine (CemSIM) at the Rensselaer Polytechnic Institute. He is currently a clinical associate professor in the Department of Surgery at the Texas A&M Health Science Center and assistant director of the Center for Evidence Based Simulation (CEBS) at the Baylor University Medical Center in Dallas, TX. His interests are in haptic device design, surgical robotics, surgical simulation, and human computer interaction. He is a member of the IEEE and IEEE Robotics and Automation Society.
[…]